### v2

# GlotEval: A Test Suite for Massively Multilingual Evaluation of Large Language Models

Abstract
Large language models (LLMs) are advancing at an unprecedented pace globally, with regions increasingly adopting these models for applications in their primary languages. Evaluating these models in diverse linguistic environments, especially in low-resource languages, has become a major challenge for academia and industry. Existing evaluation frameworks suffer from inconsistency across different benchmarks, being disproportionately focused on English and a handful of high-resource languages, thereby overlooking the realistic performance of LLMs in multilingual and lower-resource scenarios. To address this critical challenge of fragmented and inconsistent multilingual evaluation, we introduce GlotEval, a unified and lightweight framework that systematically integrates 27 benchmarks under a standardized ISO 639-3 language identifier system, allowing for seamless incorporation of new benchmarks. Supporting nine key tasks (machine translation, text classification, summarization, open-ended generation, reading comprehension, sequence labeling, intrinsic evaluation, instruction following and reasoning), spanning over dozens to hundreds of languages, GlotEval uniquely enables language-specific, cross-benchmark analysis and non-English-centric evaluations at a scale previously less practical for many researchers. This enables a precise diagnosis of model strengths and weaknesses in diverse linguistic contexts. A multilingual translation case study demonstrates GlotEval’s applicability for multilingual and language-specific evaluations.
GlotEval: github.com/MaLA-LM/GlotEval
GlotEval: A Test Suite for Massively Multilingual Evaluation of Large Language Models
[…]
1 Introduction
In recent years, driven by rapid progress in natural language processing and deep learning, large language models (LLMs) such as GPT-4 (OpenAI, 2023) and DeepSeek-R1 (DeepSeek-AI et al., 2025) have shown remarkable reasoning and generation capabilities across multiple languages and tasks. Although these models approach or surpass expert-level performance in certain high-resource languages (e.g., English), they often exhibit substantial performance fluctuations in other linguistic environments (Zhang et al., 2024). This discrepancy partially arises from the imbalance and scarcity of training data of low-resource languages, and partially from the limited multilingual coverage of current evaluation frameworks: many were originally designed for English or a few widely spoken languages, making it difficult to extend them efficiently to more diverse linguistic tasks or to adapt custom prompts and configurations for each language. Meanwhile, as LLMs proliferate worldwide and different regions rely on their respective local languages, large-scale (massively) multilingual evaluation involving numerous low-resource languages has emerged as a critical research direction.
Recent developments in LLM evaluation toolkits such as EleutherAI’s LM Evaluation Harness (Gao et al., 2023) and UltraEval (He et al., 2024) have facilitated automatic evaluation. However, significant gaps persist in language coverage, task diversity, and evaluation flexibility (Chang et al., 2024), especially in evaluating multilingual LLMs in a massively multilingual scenario. To address these issues, we present GlotEval, an evaluation framework designed to provide systematic support for a broad range of languages, with a strong focus on low-resource ones. Building on the core processes of LLM evaluation—data preparation, model inference, post-processing, and metric computation—GlotEval introduces three novel features.
  1. 1.
    Consistent Cross-benchmark Multilingual Evaluation. We integrate 27 existing multilingual benchmarks into a unified pipeline, by standardizing all ISO 639-3 language codes in the different benchmarks, [^1] https://iso639-3.sil.org/about which is an accepted standard with a good coverage of the world’s languages. By aligning benchmark language identifiers with ISO 639-3 codes, we enable evaluations for specific languages or language groups (e.g., Bantu, Dravidian, or Uralic languages), allowing the framework to automatically search among integrated benchmarks to find matching test sets. This mapping also makes it easier to incorporate new large-scale benchmarks that target mid- or low-resource languages, ensuring flexibility for future expansions.
  2. 2.
    Language-Specific Prompt Templates. Users can configure prompts for each language individually, thereby enabling more precise assessments of a model’s instruction-following ability across diverse linguistic settings. All templates are maintained in a centralized prompt library that supports multilingual benchmarks, allowing easy customization as needed. In this way, each task within a benchmark can be run potentially using prompts in the task’s original language, rather than defaulting to English prompts. To simplify cross-lingual adaptation, we also implemented Microsoft Translator integration that automatically propagates user-defined prompt templates from one single language to 130+ supported languages. [^2] https://learn.microsoft.com/en-us/azure/ai-services/translator/language-support
  3. 3.
    Non-English-Centered Machine Translation Evaluation. GlotEval is designed to break away from the traditional English-centric paradigm. Thanks to translation benchmarks featuring fully or partially multi-aligned datasets, GlotEval enables non-English-centered translation evaluations by allowing any supported language to serve as the pivot: users simply update the pivot language in the configuration to assess “any-to-pivot” / “pivot-to-any” translation directions. This flexibility ensures that GlotEval breaks from the traditional “English $\leftrightarrow$ other language” paradigm and adapts seamlessly to diverse, potentially low-resource, language pairs.
By bringing all these capabilities together in a cohesive framework, GlotEval aims to facilitate large-scale, in-depth evaluations of multilingual LLMs across both widely spoken and underrepresented languages, ultimately driving forward more inclusive LLM evaluation. Thus, GlotEval’s primary contribution is not the collection of new tasks, but the synergistic integration and standardization of existing benchmarks, which can be a robust tool for researchers and developers conducting massively multilingual LLM evaluation.
2 Related Work
Several evaluation toolkits and benchmarks have been developed to systematically assess LLMs. EleutherAI’s LM Evaluation Harness (Gao et al., 2023) is a widely adopted framework covering over 60 tasks, including multilingual datasets such as XNLI (15 languages) and Belebele (122 languages). UltraEval (He et al., 2024) improves modularity and supports FLORES-200 for multilingual translation. OpenAI Evals provides a highly flexible, community-driven framework, [^3] https://github.com/openai/evals and OpenCompass (Contributors, 2023) offers a comprehensive platform with broad support for datasets and models. MEGA (Ahuja et al., 2023) evaluates generative LLMs across diverse languages, with a focus on standard NLP benchmarks. LightEval (Fourrier et al., 2023) developed a flexible LLM evaluation framework that supports different backends.
Despite these advancements, significant gaps remain in language coverage, task diversity, and evaluation flexibility (Chang et al., 2024). Specifically, most toolkits rely on static task definitions and rarely adopt standardized language identifiers across benchmarks, making it difficult to conduct language-specific evaluations in a cross-benchmark setting. As a result, evaluations for a given language (group) must often be performed in isolation for each benchmark, limiting scalability and linguistic granularity. Furthermore, support for language-specific prompt customization is limited—most toolkits default to using English prompts regardless of the task language, which failed to take both goals of languages in multilingual evaluation, i.e., task performance versus language understanding, into consideration (Poelman and de Lhoneux, 2024).
3 GlotEval
3.1 Benchmarks, Languages and Metrics
Task Benchmark Languages Domain Open Source Metrics
Text Classification Taxi-1500 Ma et al. (2024) 1507 Bible text Yes (GitHub) Acc., F1
SIB-200 Adelani et al. (2024) 205 News topics Yes (HF, GitHub) Acc., F1
Token Classification WikiANN Pan et al. (2017) 282 Wikipedia NER Yes (HF) F1
UD treebank v2.15 de Marneffe et al. (2021) 148 POS tagging Yes (UD website) F1
Machine Translation FLORES-200 NLLB Team et al. (2022) 200+ General web Yes (HF) BLEU, ChrF++, COMET
FLORES+ 212 Gen. web, low-resource focus Yes (HF) BLEU, ChrF++, COMET
NTREX-128 Federmann et al. (2022) 128 News Yes (GitHub) BLEU, ChrF++, COMET
AmericasNLP de Gibert et al. (2025) 14 Short sentences, court proceedings, books. Yes (GitHub) BLEU, ChrF++
TICO-19 Anastasopoulos et al. (2020) 37 COVID-19 medical Yes (GitHub, OPUS) BLEU, ChrF++
IN22 Gala et al. (2023) 23 Indian langs., news+conv. Yes (GitHub) BLEU, ChrF++
NTEU Bié et al. (2020) 24 EU formal (gov) Partial (Upon request) BLEU, ChrF++
MAFAND Adelani et al. (2022) 22 News Yes (GitHub) BLEU, ChrF++
Tatoeba Challenge v2023 Tiedemann (2020) 500+ Mixed short sents. Yes (GitHub) BLEU, ChrF
OpenSubtitles v2024 Lison and Tiedemann (2016) 93 Subtitles Yes (GitHub) BLEU, ChrF
MMHB Tan et al. (2024) 9 Multilingual bias detection Yes (GitHub) ChrF with gender
Open-Ended Generation Aya Singh et al. (2024b) 119 Instruction-following Yes (HF) self-BLEU
PolyWrite Ji et al. (2024) 240 Creative writing Yes (HF) self-BLEU
Intrinsic Evaluation PBC Mayer and Cysouw (2014) 372+ Bible text Partial (Upon request) NLL
MaLA (Ji et al., 2024) 546 General web Yes (HF) NLL
Comprehension MMMLU Hendrycks et al. (2021) 14+ General knowledge QA Yes (HF) Acc.
Global-MMLU Singh et al. (2024a) 42 Culture-aware QA Yes (HF) Acc.
Summarization XLSum Hasan et al. (2021) 44 News Yes (HF, GitHub) ROUGE
MassiveSumm Long (Varab and Schluter, 2021) 55 News Yes (HF) ROUGE
MassiveSumm Short (Varab and Schluter, 2021) 88 News Yes (HF) ROUGE
Instruction Following BenchMAX Rule-based (Huang et al., 2025) 17 Verifiable instructions Yes (HF) Instruction-level Acc. etc.
Reasoning BenchMAX Math (Huang et al., 2025) 17 Grade School Math Yes (HF) Accuracy
BenchMAX Science (Huang et al., 2025) 17 Graduate-level Scientific QA Yes (HF) Accuracy
Table 1: Overview of multilingual LLM evaluation benchmarks, with typical metrics used in each.
As shown in Table 1, GlotEval integrates publicly available multilingual benchmark datasets, covering machine translation, text classification, summarization, open-ended generation, reading comprehension, sequence labeling, intrinsic evaluation, instruction following and reasoning, spanning a wide range of languages from high-resource to low-resource. In total, GlotEval comprises 9 tasks and 27 benchmarks, evaluates in over 1500 languages, and utilizes diverse metrics. Refer to Appendix A for more details of supported benchmark datasets.
3.2 Workflow
Figure 1: Workflow of GlotEval
As shown in Figure 1, the workflow of GlotEval proceeds from specifying which benchmarks and languages to use, to producing final metrics and visualization.
First, users specify their choices and through command-line arguments. Users can specify the language(s) and the benchmark task(s) to evaluate. Besides, as for prompting strategy choice, GlotEval supports two prompting strategies: Setting prompting strategy as single along with a chosen prompt language (e.g., eng_Latn) applies the same prompt in one single language for every dataset in one benchmark. This is useful for controlling variables or using a single reference prompt style; Setting prompting strategy as multi makes GlotEval search for a language-specific template in the prompt library, which corresponds to the tested language, falling back to English if not found. Especially in machine translation tasks, the source language typically determines the prompt’s language by default. Further, users can freely modify or expand the prompt library with a built-in multilingual prompt builder.
Upon selecting the desired benchmarks, languages, and prompt strategy, the user triggers GlotEval’s data loader to automatically locate each dataset and load the relevant language subsets. It then initializes the appropriate model backend depending on the task type. Specifically, for non-generative tasks, we employ the HuggingFace Transformers backend (Wolf et al., 2020) to ensure more efficient use of computational resources. For generation tasks, such as machine translation, summarization, and open-ended text generation, we prioritize the vLLM backend (Kwon et al., 2023) to ensure high throughput, while retaining the HF Transformers generation interface for compatibility purposes.
After model inference is completed, GlotEval automatically computes evaluation metrics according to the task-specific settings listed in Table 1. Optionally, as mentioned before, by appending --store_details, users can export each sample’s prompt, model output, reference, and corresponding scores to a JSONL file, which allows researchers to work outside the framework and conduct custom error analysis and result visualization. This ensures that our framework is not just an evaluation executor, but also a starting point for more fine-grained analysis.
3.3 A Deeper Look at Benchmark Data Loader
Figure 2: GlotEval benchmark data loader
Figure 2 illustrates the overall workflow within GlotEval’s data loading and prompt preparation pipeline. At its core, GlotEval aligns language identifiers from various benchmarks to a unified ISO 639-3_Script format. Once the alignment is complete, the standardized language codes serve as the central connection for downstream operations. When the user queries GlotEval with a target language (e.g., zho for Chinese or spa for Spanish), the system consults the language-to-code dictionary and retrieves all benchmark-specific subsets whose original language codes map to the same standardized form. These subsets are then included in the evaluation process. Moreover, if a language-specific prompting strategy is selected, GlotEval uses the same aligned codes to retrieve the appropriate prompt templates from the multilingual prompt library. For example, as shown in Figure 2, querying zho and spa will automatically select the corresponding benchmark subsets and load their respective prompts (zho_Hans, spa_Latn) for evaluation. This workflow builds on both the language code alignment mechanism and the multilingual prompt builder described in the following sections.
Language Code Alignment to ISO 639-3
Different benchmarks often use inconsistent codes for the same language (e.g., zh, zho, cmn, Chinese, Mandarin-CN etc. for Mandarin Chinese). Before reading benchmark datasets via dedicated data loaders, GlotEval unifies these language identifiers used across different benchmarks, to enable cross-benchmark language-specific evaluation and prompting. Figure 3(a) visualizes this process.
Specifically, we process each benchmark-provided language code—which may appear in the form of ISO 639-1, 639-2/B (bibliographic), 639-2/T (terminological), ISO 639-3 codes, or even language names—by utilizing the iso639-lang Python package. [^4] https://pypi.org/project/iso639-lang/ This allows us to retrieve all available mappings from the ISO 639-3 standard, including ISO 639-3 identifiers, ISO 639-2/B, 639-2/T, and ISO 639-1 codes. Using both exact and fuzzy matching strategies, we attempt to automatically identify the corresponding ISO 639-3 code for each language. A report is generated that documents, for each benchmark language, whether the match was exact or fuzzy, and whether it corresponds to an individual language or a macrolanguage in the ISO 639-3 standard.
We further identify the script used by each dataset, using GlotScript (Kargaran et al., 2024) to detect the dominant script. [^5] https://pypi.org/project/GlotScript/ Here we assume each dataset is primarily in one script. We randomly select up to 100 lines and attempt script recognition into ISO 15924 script code. This ensures each dataset obtains a <language>_<script> label, such as eng_Latn. The final ISO 639-3 code, along with the script code, is stored as the value in a language-to-code dictionary within the benchmark’s configuration file. Hence, each language + script combination is standardized in GlotEval for consistent usage across benchmarks.
Multilingual Prompt Builder
We constructed a dedicated command-line prompt builder to automatically prepare or adapt prompt templates for multilingual tasks. Figure 3(b) visualizes this process. In particular, the builder leverages Microsoft Translator to convert an instruction and/or few-shot prompt template from a given source language into 130+ target languages, while ensuring that placeholders (e.g., {src_text}) remain intact during translation. These newly created multilingual prompts, are stored in the updated prompt library. As a result, each dataset’s prompts are aligned with the same <language>_<script> language taxonomy, enabling consistent, language-specific evaluation.
Note that the automatic translation of prompts is intended as a convenience feature to support rapid, large-scale multilingual evaluation. While translation quality may vary—particularly for low-resource languages—this approach offers a practical starting point for exploratory analysis with language-specific prompts at scale. The framework remains fully customizable: users are able to provide their own human-written or verified prompts in the prompt library for languages of interest.
Figure 3: Benchmark data loader components: (a) Language ID alignment process and (b) multilingual prompt generation.
4 Evaluation
4.1 Efficiency Analysis
We benchmark GlotEval’s inference speed on six tasks: FLORES-200, Aya, and XLSum for generative tasks, and SIB-200, Global-MMLU, and WikiANN for non-generative tasks. All evaluations are conducted on 19 languages spanning diverse writing systems (e.g., Latin, Arabic, Cyrillic, Devanagari, Chinese, etc.). For each language, we sample 10 examples per task for evaluation. We choose Qwen2-1.5B model Yang et al. (2024) for evaluation. For generative tasks, we measure generation throughput (prefilling and decoding) with vLLM backend. For non-generative tasks, we measure classification throughput (prefilling only) using HF Transformers.
We consider two GPU environments:
  • AMD MI250X 64GB (BF16, single GPU, batch size set as 1)
  • NVIDIA A100 40GB (BF16, single GPU, batch size set as 1)
For detailed throughput performance, Appendix B shows statistics on both GPU environments. They demonstrate that in general, NVIDIA A100 consistently achieves higher throughput than AMD MI250X across both generative and non-generative tasks. Besides, this gap may also reflect the different backends between vLLM and HF Transformers. We further observe that scripts such as Devanagari or Amharic (amh_Ethi) often have lower throughput, potentially due to more complex tokenization. Lastly, summarization tasks like XLSum typically involve longer inputs and outputs than sentence-level translation tasks (e.g., FLORES-200), which increases the prefilling overhead and thus reduces the overall tokens/s.
4.2 Case Study on Multilingual Translation
Prompt Strategy Tested Translation Language Pair Prompt Template
multi
language-specific
fra → fin
French → Finnish
Traduisez la phrase suivante de Langue française en Langue finnoise
[Langue française] : {source_text_in_finnish}
[Langue finnoise] :
fin_Latn
Finnish
vie → zho-CN
Vietnamese → Chinese (Simplified)
Käännä seuraava lause Vietnamin kieli muotoon Kiinan kieli (yksinkertaistettu)
[Vietnamin kieli]: {source_text_in_vietnamese}
[Kiinan kieli (yksinkertaistettu)]:
Table 2: A demonstration of prompt templates of translation tasks in different prompt strategies.
To further illustrate GlotEval’s capabilities, we conducted a detailed case study comparing EMMA-500 (Ji et al., 2024), a large-scale multilingual language model designed to enhance multilingual performance, with the base Llama-2-7B model (Touvron et al., 2023) across various multilingual translation scenarios. This study aimed to investigate performance differences under different prompting strategies and diverse language-centric translation tasks. We designed a factorial experiment with the following variables:
  • Models: EMMA-500 vs. Llama-2-7B
  • Prompting strategies: multilingual prompting (source language-specific), Chinese prompting (zho_Hans), Finnish prompting (fin_Latn), and English prompting (eng_Latn)
[…]
A demonstration of prompt templates is shown in Table 2. For evaluation, we utilized NTREX-128, a multi-aligned benchmark containing parallel texts across 128 languages, which is supported in GlotEval. In the multilingual prompting condition, we used built-in prompt builder in GlotEval, with the support of Microsoft Translator service, to automatically translate prompts into 134 languages supported by their platform. In our case study, 106 of these languages overlap with NTREX-128 languages, allowing us to test performance across this diverse language set.
Figure 4: ChrF scores for different translation directions comparing EMMA-500 and Llama-2-7B across four prompting strategies.
The results of our case study (Figure 4) clearly demonstrate EMMA-500’s performance compared to Llama-2-7B in multilingual instruction following capabilities and non-English-centric translation tasks. Specifically, EMMA-500 shows consistently higher ChrF scores across most language pairs for all six translation directions. This performance advantage is particularly pronounced when using non-English prompting strategies, highlighting EMMA-500’s enhanced ability to process and respond to instructions in diverse languages.
The experimental design was implemented using GlotEval, which facilitated the systematic manipulation of variables through simple configuration settings. By simply modifying the prompting strategy parameter and central language settings in the multi-aligned MT benchmark configuration, we are able to comprehensively assess the language models’ multilingual capabilities, including both instruction following and non-English-centric multilingual translation.
5 Conclusion and Future Work
In this work, we introduced GlotEval, a lightweight yet comprehensive framework for massively multilingual evaluation of LLMs. By supporting consistent multilingual benchmarking, incorporating language-specific prompt templates, and supporting flexible non-English-centric translation setups, GlotEval enables consistent assessments of LLMs in diverse linguistic contexts—including low-resource settings often neglected by traditional benchmarks. Our case study on multilingual machine translation with two LLMs illustrates the utility of GlotEval in revealing the strengths and weaknesses of multilingual LLMs and in identifying directions for future optimization. Overall, GlotEval aims to encourage more inclusive, transparent, and holistic evaluations of language models across a wide array of languages and tasks, thereby advancing robust multilingual NLP research.
As for future work, we plan to integrate more diverse and comprehensive multilingual benchmarks to better evaluate LLM performance. Plus, we will explore the integration of benchmarks that the synergistic combination of automatic and human evaluation; for example, this could achieved through our pilot development of a lightweight web interface that supports crowd-sourced and expert-driven evaluation to supplement the automatic evaluation. [^6] Source code and documentation are available at https://github.com/MaLA-LM/GlotEval-HumanEval and https://gloteval-humaneval.readthedocs.io
Ethical Considerations and Broader Impact
Ethical Considerations
We strive to uphold the principles outlined in the ACL Code of Ethics. While GlotEval advances multilingual evaluation, several limitations remain. Many benchmarks still lack sufficient or high-quality data for truly low-resource languages, potentially skewing performance assessments. Additionally, as noted by Joshi et al. (2025), existing datasets often inherit cultural and linguistic biases, favoring dominant dialects or standardized language forms over regional or marginalized variants. Computational costs further constrain accessibility: large-scale evaluations are resource-intensive, posing barriers for smaller research teams. More critically, reference-free metrics introduce inherent biases, as they effectively pit one generative model against another (Deutsch et al., 2022). Such metrics struggle to capture fluency, accuracy, or cultural appropriateness, particularly in low-resource contexts where human judgments are essential.
Broader Impact
GlotEval promotes equitable progress in NLP by enabling systematic evaluation of large language models (LLMs) across diverse languages. We aim to support researchers and developers in creating language technologies that serve diverse communities more effectively via a more inclusive and holistic evaluation suite.
Acknowledgments
This project is funded by the AI-DOC program hosted by Finnish Center of Artificial Intelligence (decision number VN/3137/2024-OKM-6).
The work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546], and the Digital Europe Programme under grant agreement No 101195233.
The authors wish to acknowledge CSC - IT Center for Science, Finland, the Leonardo and LUMI supercomputers, owned by the EuroHPC Joint Undertaking, for providing computational resources.
Sawal Devkota, Bhavani Sai Praneeth Varma Mantina, Ananda Sreenidhi, Mengjie Wang, and Samea Yusofi contributed to this project as part of the “Data Analysis Software Project for Natural Language” course at TU Darmstadt, under the guidance of Shaoxiong Ji. This teaching activity was funded by LOEWE Center DYNAMIC as part of the Hessian program for the promotion of cutting-edge research LOEWE under the grant number of LOEWE1/16/519/03/09.001(0009)/98.
Appendix A Benchmark Settings
A.1 Intrinsic Evaluation
Given the input $X=(x_{0},x_{1},\ldots,x_{n_{t}})$, the negative log-likelihood (NLL) is defined as:
$\text{NLL}=-\sum_{i=1}^{n_{t}}\log p_{\theta}(x_{i}|x_{<i})$ (1)
while perplexity (PPL) is computed as:
$\text{PPL}=\exp\{-\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\log p_{\theta}(x_{i}|x_{<i})\}$ (2)
Intuitively, PPL evaluates a model’s ability to predict tokens in a given corpus, with lower values indicating better performance. In contrast, NLL measures the overall likelihood of the corpus under the model. Notably, due to its length normalization, PPL is directly influenced by the tokenization scheme, whereas NLL remains unaffected. Therefore, we use NLL for model comparisons to ensure consistency across models with different tokenization methods.
We compute NLL by concatenating the input sentences and applying a strided sliding window of size 1024.
A.2 Machine Translation
FLORES+
This work builds upon previous efforts on multilingual machine translation and evaluation datasets NLLB Team et al. (2024); Goyal et al. (2022); Guzmán et al. (2019); Doumbouya et al. (2023); AI4Bharat et al. (2023); Perez-Ortiz et al. (2024); Abdulmumin et al. (2024); Ali et al. (2024); Kuzhuget et al. (2024); Yu et al. (2024); Mamasaidov and Shopulatov (2024); Gordeev et al. (2024).
AmericasNLP
Only the development set is used, as the test set is not disclosed. Note that this dataset is aligned with Spanish, but not English.
Tatoeba (v2023-09-26)
We keep only test sets with over 1,000 sentences.
BLEU
In our experiments, BLEU scores are computed via SacreBLEU (Post, 2018) with the flores200 tokenizer to quantify translation quality. The BLEU signature is:
nrefs:1 | case:mixed | eff:no | tok:flores200 | smooth:exp | version:2.4.2
COMET
Users can specify the customized model in the configuration file. The default model is Unbabel/wmt22-comet-da.
ChrF with Gender
ChrF with gender is an evaluation metric that calculates the standard chrF score separately for sentences marked with different grammatical genders (masculine and feminine). By comparing these scores, one can assess whether a translation system favors one gender form over the other, thereby revealing potential gender bias in its outputs.
A.3 Text Classification
In classification tasks, the model predicts by ranking logits for each category; candidate labels are tokenized, and the label corresponding to the token with the highest probability is selected.
Appendix B Throughput Statistics
GlotEval provides a uniform pipeline for measuring both decoding-heavy and classification-style tasks across different languages, scripts, and hardware setups. According to efficiency analysis conducted in section 4.1, table 3 and 4 show throughput results on both NVIDIA A100 40GB and AMD MI250X 64GB GPU environments.
Language FLORES-200(Eng-X) Aya XLSum SIB-200 Global-MMLU WikiANN
(3-shot) (0-shot) (0-shot) (3-shot) (0-shot) (3-shot)
French (fra_Latn) 854 / 0.88 = 969.55 447 / 0.77 = 583.55 67 / 0.09 = 720.32 10 / 0.53 = 18.88 10 / 0.27 = 36.17 70 / 2.36 = 29.60
Swahili (swa_Latn) 1174 / 0.92 = 1274.74 812 / 0.80 = 1020.78 150 / 0.56 = 268.32 10 / 0.56 = 17.71 10 / 0.31 = 32.65 61 / 1.97 = 30.91
Vietnamese (vie_Latn) 1206 / 0.92 = 1304.01 443 / 0.76 = 581.40 172 / 0.74 = 233.62 10 / 0.48 = 20.99 10 / 0.26 = 37.87 74 / 2.41 = 30.66
Indonesian (ind_Latn) 776 / 0.87 = 893.11 259 / 0.75 = 346.56 308 / 0.75 = 411.16 10 / 0.53 = 18.91 10 / 0.28 = 35.65 54 / 2.02 = 26.79
Latin Scri. 4010 / 3.59 = 1116.99 1961 / 3.08 = 636.69 697 / 2.14 = 325.70 40 / 2.10 = 19.05 40 / 1.12 = 35.71 259 / 8.76 = 29.57
Kyrgyz (kir_Cyrl) 1174 / 0.93 = 1259.10 436 / 0.76 = 573.19 324 / 0.75 = 429.72 10 / 0.72 = 13.95 10 / 0.35 = 28.98 72 / 4.20 = 17.16
Russian(rus_Cyrl) 1280 / 1.86 = 688.45 551 / 0.77 = 712.23 339 / 0.67 = 507.16 10 / 0.53 = 18.92 10 / 0.28 = 35.25 71 / 3.51 = 20.20
Serbian (srp_Cyrl) 1118 / 0.92 = 1207.56 475 / 0.76 = 621.45 342 / 0.76 = 452.07 10 / 0.62 = 16.25 10 / 0.30 = 33.14 48 / 1.94 = 24.76
Ukrainian (ukr_Cyrl) 1083 / 0.91 = 1191.05 404 / 0.76 = 532.91 43 / 0.09 = 470.72 10 / 0.68 = 14.65 10 / 0.31 = 31.68 132 / 7.43 = 17.78
Cyrillic Scri. 4655 / 4.62 = 1007.58 1866 / 3.05 = 611.80 1048 / 2.27 = 461.67 40 / 2.55 = 15.69 40 / 1.24 = 32.26 323 / 17.08 = 18.91
Arabic (arb_Arab) 852 / 0.87 = 974.46 74 / 0.41 = 181.59 228 / 1.62 = 140.36 10 / 0.53 = 18.85 10 / 0.28 = 36.32 76 / 2.75 = 27.65
Persian (fas_Arab) 852 / 0.89 = 958.99 264 / 0.75 = 353.62 333 / 0.76 = 440.70 10 / 0.68 = 14.63 10 / 0.31 = 31.74 54 / 12.48 = 4.32
Arabic Scri. 1704 / 1.76 = 968.18 338 / 1.16 = 291.38 561 / 2.38 = 235.71 20 / 1.21 = 16.53 20 / 0.59 = 33.90 130 / 15.23 = 8.54
Bengali (ben_Beng) 1143 / 0.96 = 1190.74 973 / 0.81 = 1195.97 260 / 0.71 = 366.53 10 / 1.26 = 7.91 10 / 0.45 = 21.99 39 / 2.13 = 18.32
Hindi (hin_Deva) 1167 / 0.96 = 1210.17 960 / 0.81 = 1182.68 223 / 0.75 = 296.00 10 / 1.10 = 9.07 10 / 0.39 = 25.62 52 / 2.50 = 20.79
Nepali (npi_Deva) 1250 / 1.01 = 1247.45 803 / 0.80 = 1009.63 231 / 0.60 = 384.25 10 / 1.02 = 9.78 10 / 0.41 = 24.57 69 / 3.98 = 17.32
Devanagari 3560 / 2.93 = 1215.02 2736 / 2.42 = 1130.58 714 / 2.06 = 346.60 30 / 3.38 = 8.88 30 / 1.25 = 24.00 160 / 8.61 = 18.58
Sinhala (sin_Sinh) 1280 / 1.04 = 1226.21 1280 / 0.86 = 1485.77 103 / 0.17 = 601.67 10 / 1.57 = 6.38 10 / 0.52 =19.38 69 / 5.43 = 12.70
Telugu (tel_Telu) 1208 / 1.02 = 1188.70 559 / 0.80 = 697.54 74 / 0.14 = 537.25 10 / 1.57 = 6.38 10 / 0.55 = 18.21 71 / 8.01 = 8.86
Amharic (amh_Ethi) 1280 / 1.00 = 1278.40 1280 / 0.85 = 1498.47 65 / 0.09 = 700.75 10 / 1.00 = 9.95 10 / 7.37 = 1.36 53 / 10.31 = 5.14
Japanese (jpn_Jpan) 714 / 0.87 = 820.25 152 / 0.21 = 707.20 274 / 0.75 = 365.11 10 / 0.48 = 21.01 10 / 0.28 = 35.99 389 / 33.70 = 11.54
Korean (kor_Hang) 1016 / 0.90 = 1129.38 284 / 0.76 = 374.84 59 / 0.12 = 493.29 10 / 0.54 = 18.58 10 / 0.27 = 36.78 91 / 5.20 = 17.50
Chinese (zho_Hans) 676 / 0.87 = 780.69 403 / 0.62 = 651.94 59 / 0.12 = 491.30 10 / 0.41 = 24.11 10 / 0.26 =37.60 419 / 42.26 = 9.91
Table 3: Throughput with NVIDIA A100 40GB GPU. Each cell contains: $\frac{\#\text{generated tokens}}{\text{wall time (seconds)}}=\text{average tokens/s}$.
Language FLORES-200(Eng-X) Aya XLSum SIB-200 Global-MMLU WikiANN
(3-shot) (0-shot) (0-shot) (3-shot) (0-shot) (3-shot)
French (fra_Latn) 800 / 1.53 = 524.33 409 / 1.34 = 304.24 164 / 1.01 = 161.69 10 / 29.00 = 0.34 10 / 39.30 = 0.25 70 / 38.18 = 1.83
Swahili (swa_Latn) 1039 / 1.55 = 670.79 136 / 0.43 = 317.94 226 / 0.93 = 244.26 10 / 26.71 = 0.37 10 / 38.61 = 0.26 61 / 38.21 = 1.60
Vietnamese (vie_Latn) 932 / 1.53 = 608.26 675 / 1.39 = 485.18 58 / 0.15 = 379.43 10 / 31.58 = 0.32 10 / 39.46 = 0.25 74 / 38.13 = 1.94
Indonesian (ind_Latn) 1076 / 1.52 = 706.44 779 / 1.40 = 555.64 262 / 1.29 = 203.48 10 / 29.33 = 0.34 10 / 39.14 = 0.26 54 / 37.16 = 1.45
Latin Scri. 3847 / 6.13 = 627.57 1999 / 4.56 = 438.38 710 / 3.38 = 210.06 40 / 29.20 = 1.37 40 / 39.22 = 1.02 259 / 37.98 = 6.82
Kyrgyz (kir_Cyrl) 1051 / 1.57 = 669.63 344 / 1.32 = 261.10 444 / 1.36 = 325.96 10 / 17.18 = 0.58 10 / 37.68 = 0.27 72 / 23.48 = 3.07
Russian(rus_Cyrl) 1280 / 1.86 = 686.47 442 / 1.37 = 322.04 243 / 1.03 = 234.98 10 / 29.37 = 0.34 10 / 39.04 = 0.26 71 / 30.55 = 2.32
Serbian (srp_Cyrl) 1210 / 1.58 = 767.56 560 / 1.38 = 406.04 261 / 1.17 = 222.39 10 / 19.57 = 0.51 10 / 38.61 = 0.26 48 / 36.64 = 1.31
Ukrainian (ukr_Cyrl) 939 / 1.55 = 607.67 378 / 1.33 = 284.88 103 / 0.42 = 244.48 10 / 17.34 = 0.58 10 / 38.31 = 0.26 132 / 23.54 = 5.61
Cyrillic Scri. 4480 / 6.56 = 682.93 1724 / 5.40 = 319.26 1051 / 3.98 = 264.07 40 / 19.90 = 2.01 40 / 38.10 = 1.05 323 / 26.24 = 12.31
Arabic (arb_Arab) 919 / 1.54 = 595.36 160 / 1.24 = 129.25 83 / 0.26 = 318.11 10 / 29.06 = 0.34 10 / 39.23 = 0.25 76 / 37.56 = 2.02
Persian (fas_Arab) 929 / 1.55 = 600.20 16 / 0.12 = 131.16 184 / 1.21 = 152.61 10 / 17.43 = 0.57 10 / 38.25 = 0.26 54 / 13.24 = 4.07
Arabic Scri. 1848 / 3.09 = 598.06 176 / 1.36 = 129.41 267 / 1.47 = 181.63 20 / 21.98 = 0.91 20 / 39.22 = 0.51 130 / 21.35 = 6.09
Bengali (ben_Beng) 1130 / 1.62 = 698.59 1026 / 1.40 = 734.29 178 / 1.21 = 147.66 10 / 11.17 = 0.90 10 / 27.49 = 0.36 39 / 28.34 = 1.38
Hindi (hin_Deva) 1160 / 1.62 = 714.58 650 / 1.41 = 462.11 186 / 1.21 = 154.24 10 / 12.17 = 0.82 10 / 34.96 = 0.29 52 / 31.38 = 1.66
Nepali (npi_Deva) 1280 / 1.66 = 768.85 1126 / 1.40 = 805.59 275 / 1.10 = 250.46 10 / 13.00 = 0.77 10 / 34.68 = 0.29 69 / 2.77 = 24.87
Devanagari 3570 / 4.90 = 728.57 2802 / 4.21 = 665.56 639 / 3.52 = 181.53 30 / 12.05 = 2.49 30 / 31.91 = 0.94 160 / 5.73 = 27.91
Sinhala (sin_Sinh) 1280 / 1.76 = 727.47 1223 / 1.42 = 858.97 140 / 0.93 = 151.08 10 / 9.15 = 1.09 10 / 25.60 = 0.39 69 / 15.70 = 4.40
Telugu (tel_Telu) 1280 / 1.73 = 737.77 507 / 1.45 = 348.78 198 / 0.92 = 214.92 10 / 9.14 = 1.09 10 / 24.87 = 0.40 71 / 11.99 = 5.92
Amharic (amh_Ethi) 1280 / 1.66 = 772.22 1153 / 1.40 = 821.85 211 / 1.12 = 189.18 10 / 13.04 = 0.77 10 / 34.30 = 0.29 53 / 32.18 = 1.65
Japanese (jpn_Jpan) 690 / 1.51 = 458.09 266 / 1.27 = 209.63 250 / 1.02 = 244.87 10 / 31.85 = 0.31 10 / 39.23 = 0.25 389 / 14.98 = 25.96
Korean (kor_Hang) 973 / 1.53 = 633.96 468 / 1.38 = 340.21 204 / 1.07 = 191.09 10 / 28.74 = 0.35 10 / 39.33 = 0.25 91 / 25.01 = 3.64
Chinese (zho_Hans) 823 / 1.52 = 540.58 248 / 1.00 = 248.61 109 / 0.39 = 276.83 10 / 35.54 = 0.28 10 / 39.37 = 0.25 419 / 13.88 = 30.20
Table 4: Throughput with AMD MI250X 64GB GPU. Each cell contains: $\frac{\#\text{generated tokens}}{\text{wall time (seconds)}}=\text{average tokens/s}$.